9. A ball is dropped from a window and hits the ground with a speed of 25 ft/s. How high above the ground is the window?

1
Expert's answer
2022-03-20T18:51:21-0400

The final velocity of a ball

"v=\sqrt{2gh}"

Hence, the window is placed above the ground at the level

"h=\frac{v^2}{2g}=\rm\frac{(25\: ft/s)^2}{2\times 32.17\: ft/s^2}=9.7\: ft"